# The behavior of rich-club coefficient in scale-free networks

### v2

@@ Abstract
The rich-club phenomenon, which provides information about the association between nodes, is a useful method to study the hierarchy structure of networks. In this work, we explore the behavior of rich-club coefficient (RCC) in scale-free networks, and find that the degree-based RCC is a power function of degree centrality with power exponent $\beta$ and the betweenness-based RCC is a linear function of betweenness centrality with slope $\theta$. The findings from the New York metro network further validate our conclusions. Moreover, we calculate the value of RCC in a BA network by deleting nodes and obtain a general expression for RCC as a function of node sequence. On this basis, the solution of RCC for centrality is also obtained, which shows that the curve of RCC is determined by the centrality distribution. In order to find out what affects $\beta$ and $\theta$, we conduct numerical simulations in scale-free networks, and observe: $\beta$ and $\gamma$ (the degree distribution exponent) increase together, $\theta$ increases with $m$ (the number of new edges) and decreases to convergence as $\gamma$ increases.
@@ I Introduction
In a large network, the influence of each node is different, with some nodes holding more significant influence and acting as “hubs”. These hubs control information flow and have great impacts on transportation efficiency of the network, which means that the hubs are critical to the overall structure and function of the networkOpsahl et al. (2008). Therefore, deliberate attacks on these hubs can result in the collapse of the entire network Zhao et al. (2018). If there are more dense connections within hubs than between hubs and other nodes, that is, hubs are more likely to connect with one another, then a rich-club will be formedZhou and Mondragón (2004). This phenomenon can be observed across many types of networks, such as social networks Aghaee et al. (2020); Ansell et al. (2016); Dong et al. (2015); Vaquero and Cebrian (2013), transport networks Wei et al. (2018); Li and Cai (2007); Zhang and Ng (2021); Zhu et al. (2021), and biological networks Van Den Heuvel and Sporns (2011); Ball et al. (2014); Griffa and Van den Heuvel (2022), etc. Essentially, the rich-club phenomenon provides insight into the strong connectivity between rich nodes (hubs). The close connection between rich nodes is the key factor behind their significant influence Jiang and Zhou (2008). This facilitates the rapid dissemination of influence among them, ultimately affecting the entire networkBerahmand et al. (2018). A strong rich-club in a social network suggests that the network is not composed of disjointed and loosely connected small groups. Rather, there is a dominant “top-level” group that serves as the core structure of the networkCinelli et al. (2017). Intuitively, nodes with higher degree have more edges, and connect to each other more easily, thus forming a more densely connected club.
@@
Figure 1: The value of RCC in BA network BA$(n,m)$, where $n$ is the number of nodes and $m$ is the number of new edges introduced at each step. (a) Log-log plot of $\phi(k)$ versus degree centrality. (b) The plot of $\phi(b)$ versus betweenness centrality. As the maximum value of betweenness centrality is too large, for displaying the feature, the top one percent of nodes has been removed from the plot.
@@
The paper is organized as follows. Sec. II introduces the functional relation between degree and betweenness in scale-free networks. In Sec. III, we employ two methods to calculate RCC, and obtain the relation between RCC and node sequence. In Sec. IV, we conduct numerical simulations in scale-free networks to study the behavior of $\beta$ and $\theta$. Finally, we conclude in Sec. V.
@@ III.1 RCC and Correlation
where $E_{i,j}$ is the number of edges between nodes with degree $i$ and nodes with degree $j$, $N$ is the number of nodes, and $\left\langle k\right\rangle$ is the average degree of the network. Then we can obtain, $2E_{>k}=\textstyle\sum_{ij}E_{ij}=\textstyle\sum_{ij}N\left\langle k\right\rangle p(i,j),N_{>k}=N\textstyle\sum_{i}p(i)$. Converting the numerator and denominator of $\phi(k)$ into calculating cumulants, we obtain,
@@
Figure 3: The inverse proportional relation between RCC $\phi(i)$ and betweenness-based node sequence $i$. The value of $\phi(i)$ overlaps well with the curve $f(i)=\frac{1.76}{595-i}$, and we also show local details in the subplot.
Figure 4: Sequences of betweenness centrality in BA networks. The nodes are sorted by their values of BC. The abscissa is the sequence number, and the ordinate is the corresponding value of BC.
@@ III.2 RCC and Centrality Distribution
Our next step is to obtain RCC of a SF network through an evolutionary process. To illustrate it, let’s consider a BA network BA$(n,m)$ with $m$ new edges being added at each step. Nodes in the network are ranked by their richness in ascending order, and each node is assigned a sequence number based on its position in the ranking list. After node $i$ is removed, the network density is
@@
The variable $l_{j}$ represents the number of edges that are removed when node $i$ is cut off from the network. The density calculated using this formula serves as RCC of the network, but its calculation poses a challenge due to the varying values of $l_{j}$ with node deletion at each step. Thus, computing the accurate RCC value is a difficult task. When the richness is DC, preferential linking associates the sequence in which nodes are removed with the same sequence in which nodes join the network. As the mean of $l_{j}$ in the BA network is $m$, using $l_{j}=m$ directly will not affect our conclusion. In this case, the total number of edges $E={\textstyle\sum_{1}^{n-m}l_{j}}+m$. And RCC is given by
@@ IV The behaviors of exponent $\beta$ and slope $\theta$
After understanding how the curve of RCC is formed, our goal is to explore what affects the values of $\beta$ and $\theta$. We will study the parameters of network construction, which are the most basic data of a network. These include the number of nodes ($n$), the number of new edges ($m$), and the degree distribution exponent($\gamma$) in SF networks. In the numerical calculation, we use two SF models for comparison and verification. One is the BA network, generated through dynamic growth with adjustable parameters $n$ and $m$. For comparison, we construct a static scale-free (SSF) network based on the degree sequence $\left\{k_{n}\right\}$ following a power-law distribution Hagberg et al. (2020); Bayati et al. (2010). Unlike the BA network, the SSF network does not have a dynamic growth process, and has adjustable $\gamma$ (We set $\gamma$ in the usual range $\left(2,3\right]$). And we can also multiply the degree sequence $\left\{k_{n}\right\}$ by $m$ to change the network density. Therefore, the SSF network is adjustable with parameters $n$, $m$ and $\gamma$.
@@ V Summary and Conclusions
We have also discussed what affect the values of $\beta$ and $\theta$ in terms of network construction parameters. In order to compare the results of BA networks, we use a static scale-free network with strong randomness, whose degree distribution is randomly generated for a given exponent $\gamma$. In many simulations, the results show that $\beta$ and $\gamma$ increase together, $\theta$ increases with $m$ and decreases to convergence as $\gamma$ increases. To understand why $\beta$ and $\theta$ behave this way, we need to have a deeper understanding of centrality and explore how the change of $\gamma$ changes the network structure. Moreover, the study of joint distribution in scale-free networks can also help us further understand the behavior of RCC.
ACKNOWLEDGMENTS: This work was supported in part by the Fundamental Research Funds for the Central Universities, China (Grant No. CCNU19QN029), and the National Natural Science Foundation of China (Grant No. 61873104).

### v3

@@ Abstract
The rich-club phenomenon, which provides information about the association between nodes, is a useful method to study the hierarchy structure of networks. In this work, we explore the behavior of rich-club coefficient (RCC) in scale-free networks, and find that the degree-based RCC is a power function of degree centrality with power exponent $\beta$ and the betweenness-based RCC is a linear function of betweenness centrality with slope $\theta$. Moreover, we calculate the value of RCC in a BA network by deleting nodes and obtain a general expression for RCC as a function of node sequence. On this basis, the solution of RCC for centrality is also obtained, which shows that the curve of RCC is determined by the centrality distribution. In the numerical simulation, we observe: $\beta$ and $\gamma$ (the degree distribution exponent) increase together, $\theta$ increases with the average degree and decreases to convergence as $\gamma$ increases.
@@ I Introduction
Scale-free networks in the real world, such as social networks, transportation networks, and biological networks, often exhibit heterogeneous connectivity and structured organization, resulting in the emergence of modules, hierarchies, high-order structures and so on Fortunato and Newman (2022); De Domenico (2023); Battiston et al. (2021); Boguna et al. (2021), affecting their resilience and functionality across various domains Artime et al. (2024). The way nodes are connected also introduces distinctions among network nodes. In a large network, the influence of each node is different, with some nodes holding more significant influence and acting as “hubs”. These hubs control information flow and have great impacts on transportation efficiency of the network, which means that the hubs are critical to the overall structure and function of the network Opsahl et al. (2008). Therefore, the robustness and resilience of networks show significant responses to these hubs, and deliberate attacks on hubs can result in the collapse of the entire network Artime et al. (2024). If there are more dense connections within hubs than between hubs and other nodes, that is, hubs are more likely to connect with one another, then a rich-club will be formed Zhou and Mondragón (2004). The rich-club is also a type of mesoscale structure, and more accurately described as a core structure Gallagher et al. (2021). This phenomenon can be observed across many types of networks, such as social networks Ansell et al. (2016); Dong et al. (2015); Vaquero and Cebrian (2013), transport networks Wei et al. (2018); Li and Cai (2007); Zhang and Ng (2021); Zhu et al. (2021), and biological networks Van Den Heuvel and Sporns (2011); Ball et al. (2014), etc. Essentially, the rich-club phenomenon provides insight into the strong connectivity between rich nodes (hubs). The close connection between rich nodes is the key factor behind their significant influence Jiang and Zhou (2008). This facilitates the rapid dissemination of influence among them, ultimately affecting the entire network Berahmand et al. (2018). A strong rich-club in a social network suggests that the network is not composed of disjointed and loosely connected small groups. Rather, there is a dominant “top-level” group that serves as the core structure of the network Cinelli et al. (2017). Intuitively, nodes with higher degree have more edges, and connect to each other more easily, thus forming a more densely connected club.
@@
Figure 1: The value of RCC in BA network BA$(n,m)$, where $n$ is the number of nodes and $m$ is the number of new edges introduced at each step. Due to numerous and contiguous nodes of the same type, we could intuitively observe the linear relation in both graphs. (a) Log-log plot of $\phi(k)$ versus degree centrality. (b) The plot of $\phi(b)$ versus betweenness centrality. As the maximum value of betweenness centrality is too large, for displaying the feature, the top one percent of nodes has been removed from the plot.
@@
The paper is organized as follows. Sec. II introduces the functional relation between degree and betweenness in scale-free networks by referencing relevant literature. In Sec. III, we employ two methods to calculate RCC, and explore what factors contribute to the relation between RCC and node centrality. In Sec. IV, we observe that $\beta$ and $\theta$ exhibit good behaviors in numerical simulations of scale-free networks, and we present and discuss these findings accordingly. Finally, we conclude in Sec. V.
@@ III.1 RCC and Correlation
where $E_{i,j}$ is the number of edges between nodes with degree $i$ and nodes with degree $j$, $N$ is the number of nodes, and $\left\langle k\right\rangle$ is the average degree of the network. Then we can obtain, $2E_{>k}=\textstyle\sum_{ij}E_{ij}=\textstyle\sum_{ij}N\left\langle k\right\rangle p(i,j),N_{>k}=N\textstyle\sum_{i}p(i)$, where $p(i)$ is the degree distribution. Converting the numerator and denominator of $\phi(k)$ into calculating cumulants, we obtain,
@@
Figure 3: The inverse proportional relation between RCC $\phi(i)$ and betweenness-based node sequence $i$. We successfully validated the inverse proportion in Eq. (13) using the New York metro network. The value of $\phi(i)$ overlaps well with the curve $f(i)=\frac{1.76}{595-i}$, and we also show local details in the subplot.
Figure 4: Sequences of betweenness centrality in BA networks. The nodes are sorted by their values of BC. The abscissa is the sequence number, and the ordinate is the corresponding value of BC. We can find that the curve of BC and sequence number $i$ also follows an inverse proportional function, expressed as $b_{i}\sim\frac{1}{n-i}$.
@@ III.2 RCC and Centrality Distribution
To elucidate how the relation between RCC and centrality is determined, we proceed to investigate the RCC of a SF network through an evolutionary process. Let’s consider a BA network BA$(n,m)$ with $m$ new edges being added at each step. The network has $n$ nodes and $E$ edges. Nodes in the network are ranked by their richness in ascending order, and each node is assigned a sequence number based on its position in the ranking list. Here, $i$ and $j$ are used to denote the sequence numbers of nodes. As we remove nodes in order until node $i$ is removed, the network density is
@@
The variable $l_{j}$ represents the number of edges that are removed when node $j$ is cut off from the network. The density calculated using this formula serves as RCC of the network, but its calculation poses a challenge due to the varying values of $l_{j}$ with node deletion at each step. Thus, computing the accurate RCC value is a difficult task. When the richness is DC, preferential linking associates the sequence in which nodes are removed with the same sequence in which nodes join the network. As the mean of $l_{j}$ in the BA network is $m$, using $l_{j}=m$ directly will not affect our conclusion. In this case, the total number of edges $E={\textstyle\sum_{1}^{n-m}l_{j}}+m$. And RCC is given by
@@ IV The behaviors of exponent $\beta$ and slope $\theta$
After understanding how the curve of RCC is formed, our goal is to explore what affects the values of $\beta$ and $\theta$. We will study the parameters of network construction, which are the most basic data of a network. These include the number of nodes ($n$), the number of new edges ($m$), and the degree distribution exponent($\gamma$) in SF networks. In the numerical calculation, we use two SF models for comparison and verification. One is the BA network, generated through dynamic growth with adjustable parameters $n$ and $m$. For comparison, we construct a static scale-free (SSF) network based on the degree sequence $\left\{k_{n}\right\}$ following a power-law distribution Platt (2019); Bayati et al. (2010). Unlike the BA network, the SSF network does not have a dynamic growth process, and has adjustable $\gamma$ (We set $\gamma$ in the usual range $\left(2,3\right]$). And we can multiply the degree sequence $\left\{k_{n}\right\}$ by $m$ to change the network density. It also increases the average degree by a factor of $m$. Therefore, the SSF network is adjustable with parameters $n$, $m$ and $\gamma$.
@@ V Summary and Conclusions
We have also discussed what affect the values of $\beta$ and $\theta$ in terms of network construction parameters. In order to compare the results of BA networks, we use a static scale-free network with strong randomness, whose degree distribution is randomly generated for a given exponent $\gamma$. In many simulations, the results show that $\beta$ and $\gamma$ increase together, $\theta$ increases with $m$ (related to the average degree) and decreases to convergence as $\gamma$ increases. To understand why $\beta$ and $\theta$ behave this way, we need to have a deeper understanding of centrality and explore how the change of $\gamma$ changes the network structure. Moreover, the study of joint distribution in scale-free networks can also help us further understand the behavior of RCC.
Acknowledgements This work was supported in part by the Fundamental Research Funds for the Central Universities, China (Grant No. CCNU19QN029), and the National Natural Science Foundation of China (Grant No. 61873104).
Code availability The code used for the performance comparisons can be found in the repository https://github.com/angryRPG/RCC-in-SF-network.